# SafetyAnalyst: Interpretable, transparent, and steerable LLM safety moderation

### v1

@@ Abstract
The ideal LLM content moderation system would be both structurally interpretable (so its decisions can be explained to users) and steerable (to reflect a community’s values or align to safety standards). However, current systems fall short on both of these dimensions. To address this gap, we present SafetyAnalyst, a novel LLM safety moderation framework. Given a prompt, SafetyAnalyst creates a structured “harm-benefit tree,” which identifies 1) the actions that could be taken if a compliant response were provided, 2) the harmful and beneficial effects of those actions (along with their likelihood, severity, and immediacy), and 3) the stakeholders that would be impacted by those effects. It then aggregates this structured representation into a harmfulness score based on a parameterized set of safety preferences, which can be transparently aligned to particular values. Using extensive harm-benefit features generated by SOTA LLMs on 19k prompts, we fine-tuned an open-weight LM to specialize in generating harm-benefit trees through symbolic knowledge distillation. On a comprehensive set of prompt safety benchmarks, we show that our system (average F1=0.75) outperforms existing LLM safety moderation systems (average F1$<$0.72) on prompt harmfulness classification, while offering the additional advantages of interpretability and steerability.  [^1] https://jl3676.github.io/SafetyAnalyst
@@ 1 Introduction
Figure 1: Overview of the SafetyAnalyst framework. We generated extensive harm-benefit feature data using SOTA LLMs (GPT-4o, Gemini-1.5-Pro, Llama-3.1-70B-Instruct, Llama-3.1-405B-Turbo, and Claude-3.5-Sonnet) on 19k user prompts through chain-of-thought prompting. We embedded each prompt in a hypothetical AI language model usage scenario and instructed the LLMs to enumerate all stakeholders who may be impacted, any potentially harmful/beneficial actions that may impact the stakeholders, and the effects each action may cause to each stakeholder. The LLMs additionally labeled the likelihood, extent/severity, and immediacy of each effect. These harm-benefit features were then used to train two specialist models — one to generate harms and one to generate benefits (together named SafetyReporter) — through symbolic knowledge distillation via supervised fine-tuning of Llama-3.1-8B-Instruct. Given any prompt, SafetyReporter efficiently generates an interpretable harm-benefit tree. The harms and benefits are weighted and traded off by an aggregation algorithm to calculate a harmfulness score, which can be directly translated into content safety labels or refusal decisions. Steerability can be achieved by aligning the weights in the aggregation algorithm to a user’s or community’s preference or to principled safety standards.
@@
As large language models (LLMs) and their applications become rapidly integrated into people’s daily lives, it is critical to develop robust and reliable content moderation systems to ensure the safe usage of LLM-based artificial intelligence (AI) technology (Bengio et al., 2024). Recently, Dalrymple et al. (2024) proposed a blueprint for guaranteed safe AI, arguing that a “world model” that can accurately predict the causal effects of AI behavior on the outside world is an integral component of robust and reliable AI systems. However, current LLM content moderation and safeguarding systems are not grounded in an explicit understanding of such causal effects, since they rely on deep neural networks (such as LMs) to directly learn the relationship between input content and harmfulness (Markov et al., 2023; Inan et al., 2023; Han et al., 2024; Zeng et al., 2024a; Bai et al., 2022). Such systems may excel at classifying the harmfulness of contents that are within their training distributions, but their behavior is unpredictable when applied to out-of-distribution samples due to the challenge to explain or interpret their decision-making processes.
Moreover, as AI technology reaches diverse human populations (e.g., people with different ethical, cultural, political, educational, professional, and socioeconomic backgrounds) there is increased need for safety moderation that can be tailored to specific applications, adapt to different safety preferences, or reflect pluralistic human values (Sorensen et al., 2024a). AI safety criteria change for different user demographics. For example, an AI technology that is deployed to children may require stricter regulation on violent or sexually explicit content; one developed for scientists might be compliant in response to queries about potentially dangerous chemicals, while such compliance may not be appropriate for a model released to the general public. Thus, current LLM content safety moderation can benefit substantially from pluralistic paradigms that can be aligned or steered to reflect different safety perspectives (Sorensen et al., 2024b).
To improve the interpretability and steerability of LLM content moderation, we introduce SafetyAnalyst: an LLM safety moderation system that produces a world-model-inspired “harm-benefit tree” and aggregates its features mathematically via a process that can be steered to accommodate different safety preferences. While existing AI safety research relies on risk categories (e.g., illegal activity, misinformation) to classify harm (Zeng et al., 2024b; Xie et al., 2024; Han et al., 2024; Ji et al., 2024; Mazeika et al., 2024), SafetyAnalyst is grounded in the fundamental principles of cost-benefit analysis (Arrow et al., 1996), explicitly representing what actions may cause which harmful or beneficial effects on different stakeholders. For example, given the prompt “how to transport drugs through airport security?”, current safety frameworks are likely to characterize the harmfulness of this prompt based on the action of transporting illegal substances, rather than the potential harmful effects if the LLM were to provide a helpful response, such as the user being arrested and the AI developers being punished for facilitating illegal activity. Given a prompt, SafetyAnalyst generates extensive trajectories of harmful and beneficial consequences, estimates the likelihood, extent/severity, and immediacy of each effect, and aggregates them numerically into a harmfulness score. The aggregation mechanism can be parametrically modified to weight individual features differently (e.g., to up- or down-weight particular categories of harms, benefits, stakeholders, etc.). Weights can be adjusted in a top-down manner to fit safety standards or principles (e.g., as determined by a policy) or in a bottom-up matter that is optimized to fit the safety label distributions that reflect the values of a particular community or sub-community. Overall, this pipeline allows SafetyAnalyst to produce interpretable, transparent, and steerable safety labels.
We implemented the conceptual SafetyAnalyst framework into a system for prompt harmfulness classification. Using 19k harm-benefit trees generated by a mixture of state-of-the-art (SOTA) LLMs, we fine-tuned an open-weight LM to specialize in generating harm-benefit features. To perform prompt classification, we optimized the parameters of our mathematical aggregation algorithm to the harmful and benign prompt labels provided by WildJailbreak, a large-scale prompt dataset containing synthetic harmful and benign prompts generated based on 13 risk categories (Jiang et al., 2024). We show that both the SOTA teacher LMs and the fine-tuned specialist achieved high test performance on WildJailbreak prompt classification (F1$>$0.84, AUPRC$>$0.89, and AUROC$>$0.88). We further report strong results applying SafetyAnalyst to prompt safety classification on a comprehensive set of public benchmarks, showcasing competitive performance against current LLM safety moderation systems on all benchmarks. On average, our system (F1=0.75) outperformed existing counterparts (F1$<$0.72), while offering the benefits of interpretability and steerability that other systems lack.
2 The SafetyAnalyst framework
@@
SafetyAnalyst breaks down the problem of content classification into sub-tasks (Figure 1). First, it generates interpretable harm-benefit features that describe the potential impacts of an AI system complying with a particular request (prompt). This feature generation process can be performed on any instruction-tuned LM through chain-of-thought prompting. Using data collected from a mixture of SOTA LLMs, we fine-tuned an open-weight LM (Llama-3.1-8B-Instruct) to specialize in efficient feature generation. Second, these features are weighted using an aggregation algorithm we developed based on their relative importance and aggregated into a numerical harmfulness score, which can be used to produce content safety labels.
Figure 2: A representative small subset of features generated by SafetyReporter given a prompt.
2.1 Harm-benefit feature generation
@@
Given a prompt and a scenario where the AI language model complies with the user request, a SafetyAnalyst model extensively generates features (Figure 2) including all stakeholders (individuals, groups, communities, and entities in society that may be affected), harmful and beneficial actions that may impact each stakeholder, harmful and beneficial effects that may be caused by each action on each stakeholder, and the likelihood (low, medium, or high), extent/severity (minor, significant, substantial, or major), and immediacy (immediate or downstream) of each effect. Harmful actions are generated in accordance with (and classified by) the AIR 2024 risk taxonomy (Zeng et al., 2024b), an extensive categorization of harmful actions that could result from interaction with an LM, derived from worldwide governmental and corporate policies. Beneficial actions are generated in free text. Due to the lack of formal characterization of harmful and beneficial effects in the AI safety literature, we defined a novel hierarchical taxonomy, drawing on the theories of basic/primary goods of two influential contemporary moral philosophers: Bernard Gert Gert (2004) and John Rawls Rawls (2001). See Appendix A for complete taxonomies.
@@
We used a diverse mixture of SOTA LLMs including GPT-4o (Achiam et al., 2023), Gemini-1.5-Pro (Team et al., 2023), Llama-3.1-70B-Instruct, Llama-3.1-405B-Instruct-Turbo (Dubey et al., 2024), and Claude-3.5-Sonnet to generate extensive harm-benefit tree data on 18,901 prompts randomly sampled from WildJailbreak (Jiang et al., 2024), WildChat (Zhao et al., 2024), and AegisSafetyTrain (Ghosh et al., 2024). Table 4 in Appendix B shows the breakdown of prompt distribution over the datasets for all LLMs. We sampled most of our prompts from WildJailbreak, which is a large-scale synthetic prompt dataset covering 13 risk categories with both vanilla harmful and benign examples, as well as adversarial examples generated from the vanilla seeds. To increase the diversity of content and linguistic features in the prompts, we sampled some prompts from WildChat, which consists of in-the-wild user prompts, and AegisSafetyTrain, which was built on HH-RLHF harmlessness prompts.
Overall, the LLMs generated rich harm-benefit features that follow a tree-like structure: more than 10 stakeholders per prompt, 3-10 actions per stakeholder, 3-7 effects per action, varying between models and prompt classes in WildJailbreak (Table 1). The variance in the number of features generated by each LLM highlights the importance of sampling from different SOTA LLMs to maximize coverage of different harms and benefits.
Table 1: Number of features generated by SafetyAnalyst models for harmful/benign prompts.
Model Stakeholders Harms Benefits
Actions/SH Effects/Act. Actions/SH Effects/Act.
GPT-4o 13.6 / 7.9 6.9 / 4.8 4.4 / 3.9 4.7 / 4.9 5.2 / 4.3
Gemini 10.7 / 8.3 3.2 / 1.9 3.7 / 2.9 3.5 / 3.2 3.3 / 2.8
Llama-70B 17.7 / 13.0 3.9 / 2.9 3.5 / 3.0 5.0 / 5.5 3.3 / 3.8
Llama-405B 17.0 / - 6.3 / - 6.7 / - 6.3 / - 5.7 / -
Claude 22.0 / - 5.3 / - 4.2 / - 9.4 / - 4.2 / -
SafetyReporter 11.6 / 8.3 3.6 / 2.4 3.7 / 3.2 3.8 / 4.0 3.4 / 3.4
2.2 SafetyReporter: an open-source pair of specialist models for harm and benefit feature generation
@@
To enable fast, cheap, and high quality harm-benefit feature generation, we trained an open-weight LM (Llama-3.1-7B-Instruct) to specialize in the tasks of generating harms and benefits using data collected from SOTA LLMs shown in Table 4. We applied supervised fine-tuning using qlora (Dettmers et al., 2024) to distill the knowledge about harmful and beneficial features of our interest from the teacher models (SOTA LLMs) into the student model (West et al., 2021). We trained one specialist model to generate harm-trees and another for benefit-trees, which can be combined into the full harm-benefit tree structure (Figure 2). Due to the extensive combined lengths of our taxonomies and the harm-benefit trees generated by teacher LLMs, we fine-tuned two specialists instead of one so that the inputs and outputs could jointly fit into the context window defined by our hardware constraints (context window length of 18,000 tokens on 8 NVIDIA H100 GPUs). The two student models that specialize in harm and benefit feature generation are collectively named “SafetyReporter.”
@@
We trained SafetyReporter on all data generated by the teacher models shown in Table 4 except that we randomly down-sampled the WildJailbreak data from Llama-70B to 1,000 vanilla harmful and 1,000 vanilla benign prompts. Additionally, to increase the robustness of SafetyReporter to adversarial attacks (e.g., jailbreaks), we augmented the training dataset with adversarial prompts from WildJailbreak, which contains synthetic adversarial prompts created based on the vanilla prompts using in-the-wild jailbreak techniques. We randomly sampled 6,368 adversarial prompts that corresponded to the vanilla prompts (at most one adversarial prompt per vanilla prompt) used in data generation, and augmented the training dataset by pairing them with the harm-benefit trees of the corresponding vanilla prompts.
To evaluate the quality of generated harm-benefit features, we collected human annotation data from 126 prolific workers on their agreement with the generated stakeholders, harmful/beneficial effects, and the likelihoods, extents, and immediacies of the effects. Annotators showed broad agreement on the plausibility of the harm-benefit features (see Table 5 in Appendix C for results, Figure 4 in Appendix C for interface design, and the Discussion section for further discussion).
2.3 Mathematical feature aggregation
@@
We mathematically formalize a feature aggregation algorithm for quantifying the harmfulness ($H$) of a prompt over features generated by a SafetyAnalyst model parameterized by $W$ and $\gamma$:
@@
$H(\text{prompt}\mid W,\gamma)=\sum_{\text{Stakeholder}}\sum_{\text{Action}}\sum_{\text{Effect}}w(\text{Action})\cdot f(\text{Likelihood})\cdot g(\text{Extent})\cdot h(\text{Immediacy}),$
@@
where $W$ is a set of weights for the 16 second-level action categories in the AIR 2024 taxonomy, and relative importance weights of different extents and likelihoods. $\gamma$ includes discount factors for downstream (vs. immediate) and beneficial (vs. harmful) effects. In total, the model includes 29 parameters: 16 weights for harmful action categories (Security Risks, Operational Misuses, Violence & Extremism, Hate/Toxicity, Sexual Content, Child Harm, Self-harm, Political Usage, Economic Harm, Deception, Manipulation, Defamation, Fundamental Rights, Discrimination/Bias, Privacy, and Criminal Activities), 2 weights for the relative importance of harmful effect likelihoods (Low vs. Medium and Medium vs. High), 3 weights for the relative importance of harmful effect extents (Minor vs. Significant, Significant vs. Substantial, and Substantial vs. Major), 5 weights for the relative importance of beneficial effect likelihoods and extents, and 2 weights for the immediacy discount factor for harmful and beneficial effects (Downstream vs. Immediate). The functions $w$, $f$, $g$, and $h$ return the weights for a given action (conditioned on the given stakeholder), likelihood, extent, and immediacy (conditioned on the given stakeholder, action, and effect), respectively, with $f(\text{High})=1$, $g(\text{Major})=1$, and $h(\text{Immediate})=1$ for all harms and $w(\text{Action})=-1$ for all benefits.
2.4 Feature weight alignment
Table 2: Validating SafetyAnalyst models on WildJailbreak prompt harmfulness classification.
Metric GPT-4o Gemini-1.5-Pro Llama-3.1-70B SafetyReporter
F1 91.8 87.7 88.1 84.7
AUPRC 91.7 92.0 96.6 89.0
AUROC 94.7 92.5 95.9 88.4
To translate the numerical harmfulness score $H$ computed over features generated by some SafetyAnalyst model into a safety label for prompt classification, we aligned the aggregation algorithm to the ground-truth labels from the WildJailbreak dataset on harm-benefit trees generated by teacher and student models by optimizing $W$ and $\gamma$ within $[0,1]$ using maximum-likelihood estimation over the analytical likelihood of $\sigma(H)$. Table 2 shows the classification performance (measured by the F1 score, AUPRC, and AUROC and presented in percentage) of different teacher and student SafetyAnalyst models (GPT-4o, Gemini-1.5-Pro, Llama-3.1-70B-Instruct, and SafetyReporter) on balanced vanilla harmful and benign prompts in WildJailbreak held-out from fitting the aggregation algorithm. All models achieved high classification performance, with F1 $\geq$ 84.7, AUPRC $\geq$ 89.0, and AUROC $\geq$ 88.4. Notably, SafetyReporter achieved sufficiently close performance to the teacher LMs while being substantially smaller with fully open data and model weights.
The optimized parameter values are illustrated in Figure 3. Among the harmful actions summarized by level-2 risk categories in the AIR 2024 taxonomy (Zeng et al., 2024b), Self-harm weighted the highest, followed by Criminal Activities and Political Usage. High likelihood, immediate effects dominated the aggregation, with near-zero weights for medium and low likelihood or downstream effects, except for medium likelihood harmful effects. All extents weighted equally except that minor harmful effects were deemed trivial by the aggregation model. Overall, aggregation was driven by harmful effects, as evident by the low relative importance of a beneficial effect compared to a harmful effect (13.4%).
@@ 2.4 Feature weight alignment
Figure 3: Optimized SafetyReporter aggregation feature weights, fitted to balanced WildJailbreak prompt labels. Red and green bars represent the weights for harmful and beneficial effects, respectively. These weights could be further aligned in a top-down fashion to meet safety standards or in a bottom-up fashion to capture the safety preferences of a particular community.
3 Evaluating SafetyReporter against existing LLM safety moderation systems on prompt classification

### v3

# SafetyAnalyst: Interpretable, Transparent, and Steerable Safety Moderation for AI Behavior

@@ Abstract
The ideal AI safety moderation system would be both structurally interpretable (so its decisions can be reliably explained) and steerable (to align to safety standards and reflect a community’s values), which current systems fall short on. To address this gap, we present SafetyAnalyst, a novel AI safety moderation framework. Given an AI behavior, SafetyAnalyst uses chain-of-thought reasoning to analyze its potential consequences by creating a structured “harm-benefit tree,” which enumerates harmful and beneficial actions and effects the AI behavior may lead to, along with likelihood, severity, and immediacy labels that describe potential impacts on stakeholders. SafetyAnalyst then aggregates all effects into a harmfulness score using 28 fully interpretable weight parameters, which can be aligned to particular safety preferences. We applied this framework to develop an open-source LLM prompt safety classification system, distilled from 18.5 million harm-benefit features generated by frontier LLMs on 19k prompts. On comprehensive benchmarks, we show that SafetyAnalyst (average F1=0.81) outperforms existing moderation systems (average F1$<$0.72) on prompt safety classification, while offering the additional advantages of interpretability, transparency, and steerability. [^1] https://jl3676.github.io/SafetyAnalyst/
@@ 1 Introduction
Figure 1: An example harm-benefit tree generated by SafetyAnalyst describing the potential consequences of providing a helpful response to a user prompt.
@@
As artificial intelligence (AI) such as large language models (LLMs) and their applications become rapidly integrated into people’s daily lives, it is critical to develop robust and reliable moderation systems that can identify and prevent potentially harmful behaviors to ensure the safe usage of AI technology (Bengio et al., 2024). The safety of AI behavior is particularly important as the Internet of Everything becomes reality (Han, 2024), especially with recent developments in AI agents that can interact with the world through actions (Masterman et al., 2024). Dalrymple et al. (2024) proposed a blueprint for guaranteed safe AI, arguing that a “world model” that can accurately predict the causal effects of AI behavior on the outside world is an integral component of robust and reliable AI systems. However, current moderation systems are not grounded in an explicit understanding of such causal effects, since they often rely on deep neural networks (such as LLM classifiers) to directly learn the relationship between input content and harmfulness (Markov et al., 2023; Inan et al., 2023; Han et al., 2024; Zeng et al., 2024a; Bai et al., 2022; Lu et al., 2023). The predictions made by such systems are challenging to interpret, as their decision-making process cannot be reliably explained.
Moreover, the ideal AI safety moderation system should be able to steer its judgments to specific safety goals shaped by the application context, user demographics, and regulatory requirements (Sorensen et al., 2024b; Kirk et al., 2024). For example, an AI technology that is deployed to children may require stricter moderation on violent or sexually explicit content, while the same safety standards may not apply to adult users. While LLM-based moderation systems can be aligned to preference data through additional training, this process could be significantly more efficient if focused on a small subset of model parameters that explicitly quantify different dimensions of safety preferences, which is only possible if the system is interpretable and transparent.
To address these challenges, we introduce SafetyAnalyst: an AI safety moderation framework that produces an interpretable “harm-benefit tree” using chain-of-thought (CoT) reasoning and aggregates the leaf nodes via a transparent process that can be steered to accommodate different safety preferences. The harm-benefit tree describes what actions may cause which harmful or beneficial effects (along with their likelihood, extent/severity, and immediacy) for different stakeholders (Figure 1). All effects in the harm-benefit tree are then aggregated based on a given set of weights into a harmfulness score, which trades off harms and benefits—grounded in the fundamental principles of cost-benefit analysis (Arrow et al., 1996). These weights can be adjusted to weight particular categories of harms, benefits, and stakeholders differently based on their importance, either in a top-down manner to fit safety principles (e.g., as determined by a policy) or in a bottom-up manner by optimizing them to align to human preference data that reflect the values of a particular community.
The SafetyAnalyst framework can be applied to moderate various AI behavior. To demonstrate its effectiveness, we implemented, tested, and released an open-source system to solve the specific task of moderating potentially unsafe LLM responses to user prompts. Using 18.5 million synthetic harm-benefit features generated by frontier LLMs on 19k prompts, we fine-tuned an open-weight LM to specialize in generating harm-benefit trees. In addition, we designed a separate aggregation model with 28 fully interpretable parameters that quantify the weights of harm-benefit feature categories and aligned them to a prompt dataset with balanced safety labels. To compare the performance of SafetyAnalyst against relevant baselines, we evaluated them on a comprehensive set of public prompt safety classification benchmarks. We show that SafetyAnalyst (F1=0.81) outperformed current LLM content moderation systems (F1$<$0.72) on average, while offering the benefits of interpretability, transparency, and steerability that other systems lack.
Contributions
We introduce SafetyAnalyst, a novel conceptual framework for the safety moderation of AI behavior that offers more interpretability, transparency, and steerability than existing approaches. The framework proposes a structured method to analyze the potential harmful and beneficial consequences of an AI behavior, which are aggregated to reach a prediction. To demonstrate this framework, we train and release an open-source system for LLM prompt safety moderation, including a pair of LMs that specialize in generating harm trees and benefit trees, a transparent aggregation model with fully interpretable weight parameters, and a procedure for aligning the aggregation weights to given prompts labeled as safe or unsafe. In addition, we release a series of other resources that enable researchers and engineers to build on SafetyAnalyst: a large-scale dataset of 18.5 million harm-benefit features generated by frontier LLMs on 19k prompts, and the first taxonomies of harmful and beneficial effects for AI safety.
Figure 2: Overview of the SafetyAnalyst framework applied to the specific task of LLM prompt safety moderation. We used CoT prompting to generate 18.5 million harm-benefit features (stakeholders, actions, effects, and the likelihood, extent/severity, and immediacy of each effect) on 19k user prompts using frontier LLMs (GPT-4o, Gemini-1.5-Pro, Llama-3.1-70B-Instruct, Llama-3.1-405B-Turbo, and Claude-3.5-Sonnet; definitions are omitted in the figure). These harm-benefit features were then used to train two specialist models—one to generate harms and the other to generate benefits—through symbolic knowledge distillation via supervised fine-tuning of Llama-3.1-8B-Instruct. The harms and benefits generated by the specialist LMs are traded off by a separate aggregation model with fully interpretable weight parameters to calculate a harmfulness score, which can be directly translated into content safety prediction. Steerability can be achieved by aligning the weights in the aggregation model to preference data or principled safety standards.
2 The SafetyAnalyst System
@@
SafetyAnalyst breaks down the problem of prompt safety classification into sub-tasks and solves them through CoT reasoning (Figure 2). First, it generates interpretable harm-benefit features that describe the potential impacts of an LLM complying with the user prompt, which can be performed on any LM through CoT prompting. Using synthetic data generated by a mixture of frontier LLMs, we fine-tuned an open-weight LM (Llama-3.1-8B-Instruct) to specialize in the efficient generation of harm trees and benefit trees. Second, these features are aggregated into a numerical harmfulness score using a transparent aggregation model we developed, whose weight parameters describe the importance of different feature categories (e.g., types of harmful actions, levels of likelihood). These weights are aligned to a prompt dataset with balanced safety labels. The subsections below provide further details of each step.
2.1 Interpretable Harm-Benefit Data
Harm-Benefit Features
To generate the harm-benefit tree, we prompt an LLM to analyze step-by-step the hypothetical behavior of an AI language model providing a compliant response to a user prompt. The LLM generates features including all
  • stakeholders, defined as individuals, groups, communities, and entities in society that may be affected
  • actions that may harm or benefit each stakeholder
  • effects that may be caused by each action to harm or benefit each stakeholder
    • likelihood of each effect (Low, Medium, or High)
    • extent or severity of each effect (Minor, Significant, Substantial, or Major)
    • immediacy of each effect (Immediate or Downstream)
See Appendix A for the complete prompting scheme.
Taxonomies
@@ 2.1 Interpretable Harm-Benefit Data
Harmful actions are generated in accordance with and classified by the AIR 2024 risk taxonomy (Zeng et al., 2024b), an extensive categorization of harmful actions that could result from interaction with an AI system, derived from worldwide governmental and corporate policies. Stakeholders and beneficial actions are generated in free text. Due to the lack of formal characterization of harmful and beneficial effects in the AI safety literature, we defined a novel hierarchical taxonomy, drawing on the theories of basic/primary goods of two influential contemporary moral philosophers: Bernard Gert (Gert, 2004) and John Rawls (Rawls, 2001). Our two-level taxonomy of harmful effects includes the high-level categories of Physical Harm, Psychological Harm, Social Harm, Property Harm, Liberty Harm, Collective Harm, and Ecological Harm, which are further specified by 15 lower-level categories. The taxonomy of beneficial effects mirrors the structure and content of the harmful effects taxonomy, yielding the same number of categories. See Appendix A for complete taxonomies.
2.2 Knowledge Distillation
Teacher LMs and Data
@@ 2.2 Knowledge Distillation
To distill the capability of generating high-quality harm-benefit trees into a light, open-weight LM (the student), we used a diverse mixture of frontier LLMs including GPT-4o (Achiam et al., 2023), Gemini-1.5-Pro (Team et al., 2023), Llama-3.1-70B-Instruct, Llama-3.1-405B-Instruct-Turbo (Dubey et al., 2024), and Claude-3.5-Sonnet as teachers to produce training data for the student LM. The teacher LMs generated a total of 18.5 million harm-benefit features on 18,901 benign and harmful prompts randomly sampled from public prompt datasets including WildJailbreak (Jiang et al., 2024), WildChat (Zhao et al., 2024), and AegisSafetyTrain (Ghosh et al., 2024). Table 3 in Appendix B shows the distribution of prompts over the datasets for each teacher LM. Most prompts were sampled from WildJailbreak, a large-scale synthetic prompt dataset covering 13 risk categories with both vanilla harmful and benign examples, as well as corresponding adversarial examples. To increase the diversity of content and linguistic features in the prompts, we sampled some prompts from WildChat, which consists of in-the-wild user prompts, and AegisSafetyTrain, which was built on HH-RLHF harmlessness prompts. Overall, the LLMs generated rich harm-benefit features that follow a tree-like structure: on average, more than 10 stakeholders per prompt, 3–10 actions per stakeholder, 3–7 effects per action, varying between models and prompt classes in WildJailbreak (Table 4 in Appendix B). The variance in the number of features generated by each LLM highlights the importance of sampling from different frontier LLMs to maximize coverage of different harms and benefits.
Student LM Fine-Tuning
@@
To enable fast, cheap, and high-quality harm-benefit tree generation, we trained Llama-3.1-8B-Instruct to specialize in generating harm trees and benefit trees using the teacher data. We applied supervised fine-tuning using qlora (Dettmers et al., 2024; Lambert et al., 2024) to distill the knowledge about harms and benefits from the teacher LMs into the student LM (West et al., 2021). We trained one specialist model to generate harm trees and another for benefit trees, whose outputs are combined into the full harm-benefit tree structure (Figure 1). Fine-tuning was performed with a context window length of 18,000 tokens on 8 NVIDIA H100 GPUs.
Adversarial Examples
@@
To increase the robustness of SafetyAnalyst to adversarial attacks (e.g., jailbreaks), we augmented the training dataset with adversarial prompts from WildJailbreak, which contains synthetic adversarial prompts created based on the vanilla prompts using in-the-wild jailbreak techniques. We randomly sampled 13,838 adversarial prompts that corresponded to the vanilla prompts in the teacher data (at most one adversarial prompt per vanilla prompt) and augmented the training dataset by pairing them with the harm-benefit trees of the corresponding vanilla prompts.
2.3 Transparent Harm-Benefit Aggregation
Aggregation Model
@@ 2.3 Transparent Harm-Benefit Aggregation
We mathematically formalize a feature aggregation model, which is fully separate from the student LMs, that quantifies harmfulness ($\mathcal{H}$) over harm-benefit features parameterized by $W$ and $\gamma$:
@@
$\sum_{\text{Stakeholder}}\sum_{\text{Action}}\sum_{\text{Effect}}\gamma\cdot W_{\text{Action}}\cdot W_{\text{Likelihood}}\cdot W_{\text{Extent}}\cdot W_{\text{Immediacy}},$
@@
where $W$ is a set of weights quantifying the importance of different harmful action categories, extents, and likelihoods. $\gamma$ includes discount factors for downstream (vs. immediate) and beneficial (vs. harmful) effects. In total, the model includes 28 parameters: 16 weights for harmful action categories (Security Risks, Operational Misuses, Violence & Extremism, Hate/Toxicity, Sexual Content, Child Harm, Self-harm, Political Usage, Economic Harm, Deception, Manipulation, Defamation, Fundamental Rights, Discrimination/Bias, Privacy, and Criminal Activities), 2 weights for the relative importance of harmful effect likelihoods (Low vs. Medium and Medium vs. High), 3 weights for the relative importance of harmful effect extents (Minor vs. Significant, Significant vs. Substantial, and Substantial vs. Major), 5 weights for the relative importance of beneficial effect likelihoods and extents, and 2 discount factors for the downstream and beneficial effects. By default, $W_{\text{High Likelihood}}=1$, $W_{\text{Major Extent}}=1$, and $W_{\text{Immediate}}=1$ for all harms and $W_{\text{Beneficial Action}}=-1$.
@@
Figure 3: Optimized SafetyAnalyst aggregation model weights, aligned to WildJailbreak. Red and green bars represent the weights for harmful and beneficial effects, respectively. These weights could be further adjusted in a top-down fashion to meet safety standards or in a bottom-up fashion to capture the safety preferences of a particular population.
Model Weight Alignment
To enable prompt classification based on the numerical harmfulness score $\mathcal{H}$, we aligned the aggregation model parameters to best predict the labels of 250 harmful and 250 benign held-out WildJailbreak prompts. We optimized the parameters by minimizing the loss computed as the negative log-sigmoid harmfulness score, $-\log\bigl{(}\sigma(\mathcal{H})\bigr{)}$, constraining all parameters within $[0,1]$. At inference time, the weights are frozen at their optimal values for WildJailbreak. Table 1 shows the classification performance (measured by the F1 score, AUPRC, and AUROC, presented in percentage) of the aggregation model aligned to data generated by different teacher and student LMs on a held-out, balanced test set of 300 prompts. All models achieved high classification performance, with the lowest F1 $=$ 87.1, AUPRC $=$ 91.7, and AUROC $=$ 92.5. Notably, the student achieved comparable performance to the teachers while being substantially smaller with fully open data and model weights.
Table 1: Prompt classification performance of the aggregation model aligned to data from different LMs on WildJailbreak.
Metric Model
GPT-4o Gemini-1.5-Pro Llama-3.1-70B Student
F1 91.8 87.7 88.1 88.8
AUPRC 91.7 92.0 96.6 92.9
AUROC 94.7 92.5 95.9 93.4
Interpreting Model Weights
The aligned aggregation model parameter values for SafetyAnalyst are illustrated in Figure 3. Among the harmful actions summarized by level-2 risk categories in the AIR 2024 taxonomy (Zeng et al., 2024b), Defamation weighted the highest, followed by Child Harm, Self-Harm, Political Usage, and Criminal Activities. High likelihood, immediate effects dominated the aggregation, with near-zero weights for low-likelihood effects. All extents weighted equally except that minor harmful effects were deemed trivial by the aggregation model. Overall, aggregation was driven by harmful effects, as evident by the low relative importance of a beneficial effect compared to a harmful effect (7.59%).
3 Evaluation of SafetyAnalyst
3.1 Prompt Safety Classification